OpenStudy (anonymous):

a circle is circumscribed about a square. If the sides of the sqare are 6,what is the approximate circumference of the circle?(round your answer to the tenths place)

OpenStudy (anonymous):

2x^2 = 6^2 -> x = 3 sqrt 2 Circ = 2 pi x = 6 sqrt 2pi = 26.7 to 1 dp